# SPaR: Self-Play with Tree-Search Refinement to Improve Instruction-Following in Large Language Models

### v1

@@ Abstract
Instruction-following is a fundamental capability of language models, requiring the model to recognize even the most subtle requirements in the instructions and accurately reflect them in its output. Such an ability is well-suited for and often optimized by preference learning. However, existing methods often directly sample multiple independent responses from the model when creating preference pairs. Such practice can introduce content variations irrelevant to whether the instruction is precisely followed (e.g., different expressions about the same semantic), interfering with the goal of teaching models to recognize the key differences that lead to improved instruction following. In light of this, we introduce SPaR, a self-play framework integrating tree-search self-refinement to yield valid and comparable preference pairs free from distractions. By playing against itself, an LLM employs a tree-search strategy to refine its previous responses with respect to the instruction while minimizing unnecessary variations. Our experiments show that a LLaMA3-8B model, trained over three iterations guided by SPaR, surpasses GPT-4-Turbo on the IFEval benchmark without losing general capabilities. Furthermore, SPaR demonstrates promising scalability and transferability, greatly enhancing models like GLM-4-9B and LLaMA3-70B. We also identify how inference scaling in tree search would impact model performance. Our code and data are publicly available at https://github.com/thu-coai/SPaR.
@@ 2.2.1 Prompt Creation
Complex prompts constructed without human intervention tend to be poorly diversified, as the types of constraints added are often distributed unevenly (Sun et al., 2024). Therefore, we adopt a taxonomy-based mechanism to make constraint types comprehensive and balanced. Our taxonomy for instruction-following constraints is derived from Cheng et al. (2024) and further refined to be more comprehensive.
After building the constraint taxonomy, we employ it to construct complex instruction-following tasks based on seed prompts. We sample a main constraint type and employ a strong LLM to add several other constraints to make the original prompt more complex. Moreover, we leverage the strong LLM to assess the validity of the generated prompt, ensuring that the constraints do not conflict with each other or create unreasonable scenarios with the original task. The detailed taxonomy and prompt can be found in Appendix A.
@@ 2.3.1 Negative Data Collection
For each prompt $x$, we first sample $K$ responses $\{y_{1},y_{2},\ldots,y_{K}\}$ from the actor model. This step ensures that there are enough negative responses to support subsequent learning. Then, for each prompt and response pair, we utilize the refiner to generate a judgment, which contains two parts: a label suggesting whether the response follows the instruction and an explanation about the assessment. To make this judgment more accurate, we incorporate the self-consistency mechanism (Wang et al., 2022), which is also applied in the subsequent refinement process. Specifically, we obtain multiple judgments from the refiner and determine the final label through majority voting, as detailed in Appendix D.4. After majority voting, we randomly select one judgment that matches the voted label to serve as the final judgment. This process allows us to identify challenging prompts that elicit responses that do not accurately follow the instructions, yielding tuples in the form of $(x,y_{\text{negative}},j)$, where $y_{\text{negative}}$ is the incorrect response and $j$ is its corresponding judgment.
@@ 3.1 Experiment Setup
In this work, we focus on enhancing the instruction-following abilities of LLMs in a self-play fashion. We utilize SFT to bootstrap models under 10B parameters as actor and refiner models. For the more advanced LLaMA3-70B-Instruct, we directly employ it in both roles. Following this, we perform a three-iteration self-play training using 10k prompts per iteration from our generated dataset. In each iteration, we apply DPO for the actor and RFT for the refiner. We refer to the trained LLaMA3-8B-Instruct as SPaR-8B, LLaMA3-70B-Instruct as SPaR-70B, GLM-4-9B-Chat as SPaR-9B, and Mistral-7B-Instruct as SPaR-7B. More implementation details are provided in Appendix C.
Baselines.
We compare our method with five popular self-improvement approaches, including:
@@ 3.3 Actor Evaluation Results
As shown in Appendix D.2, we assessed each iteration’s performance on general benchmarks, including GSM8k (Cobbe et al., 2021), TriviaQA (Joshi et al., 2017), MMLU (Hendrycks et al., 2020), and HumanEval (Chen et al., 2021). The results indicate that SPaR maintains or even improves general performance, particularly on GSM8k and HumanEval benchmarks, demonstrating that enhanced instruction-following capabilities support overall LLM alignment.
@@
SPaR outperforms other baselines significantly.
Figure 3 demonstrates the improvements on IFEval with each training iteration. In every iteration, SPaR outperforms other methods. Notably, even after three iterations, other methods fail to surpass the performance of SPaR’s first iteration. Generally, our method and SELF outperform self-rewarding and meta-rewarding approaches, underscoring the importance of learning from refinement and excluding the interfering factors in instruction-following tasks. Furthermore, SPaR’s superior performance compared to SELF indicates that contrastive refinement response pairs can highlight key differences, which are difficult to learn using only correct responses. Additionally, only SPaR-8B-SFT outperforms the original LLaMA3-8B-Instruct, which suggests that incorporating the judgment SFT or refinement SFT data would reduce performance, likely due to the huge task gap and reduced diversity in the data.
@@ 3.5 Ablations and Analysis
Figure 5: Comparison of decoding strategies. Model performance improves with increased inference times.
@@ Appendix C Implementation Details
Appendix D Experiment Results
D.1 Instruction-Following Evaluation Results.
@@ Appendix D Experiment Results
D.2 General Performance Evaluation
@@
D.3 Judgment Evaluation Results.
@@
D.4 Ablation Study on Judgment Capability.
@@
D.5 Ablation Study on Refinement Capability.
@@
D.6 Inference-time Scaling Comparison

### v2

@@ Abstract
Instruction-following is a fundamental capability of language models, requiring the model to recognize even the most subtle requirements in the instructions and accurately reflect them in its output. Such an ability is well-suited for and often optimized by preference learning. However, existing methods often directly sample multiple independent responses from the model when creating preference pairs. Such practice can introduce content variations irrelevant to whether the instruction is precisely followed (e.g., different expressions about the same semantic), interfering with the goal of teaching models to recognize the key differences that lead to improved instruction following. In light of this, we introduce SPaR, a self-play framework integrating tree-search self-refinement to yield valid and comparable preference pairs free from distractions. By playing against itself, an LLM employs a tree-search strategy to refine its previous responses with respect to the instruction while minimizing unnecessary variations. Our experiments show that a LLaMA3-8B model, trained over three iterations guided by SPaR, surpasses GPT-4-Turbo on the IFEval benchmark without losing general capabilities. Furthermore, SPaR demonstrates promising scalability, greatly enhancing models like GLM-4-9B and LLaMA3-70B. We also identify how inference scaling in tree search would impact model performance. Our code and data are publicly available at https://github.com/thu-coai/SPaR.
@@ 2.2.1 Prompt Creation
Complex prompts constructed without human intervention tend to be poorly diversified, as the types of constraints added are often distributed unevenly (Sun et al., 2024). Therefore, we adopt a taxonomy-based mechanism to make constraint types comprehensive and balanced. Our taxonomy for instruction-following constraints is derived from Cheng et al. (2024) and further refined to be more comprehensive. After building the constraint taxonomy, we employ it to construct complex instruction-following tasks based on seed prompts. We sample a main constraint type and employ a strong LLM to add several other constraints to make the original prompt more complex. Moreover, we leverage the strong LLM to assess the validity of the generated prompt, ensuring that the constraints do not conflict with each other or create unreasonable scenarios with the original task. The detailed taxonomy and prompt can be found in Appendix A.
@@ 2.3.1 Negative Data Collection
For each prompt $x$, we first sample $K$ responses $\{y_{1},y_{2},\ldots,y_{K}\}$ from the actor model. This step ensures that there are enough negative responses to support subsequent learning. Then, for each prompt and response pair, we utilize the refiner to generate a judgment, which contains two parts: a label suggesting whether the response follows the instruction and an explanation about the assessment. To make this judgment more accurate, we incorporate the self-consistency mechanism (Wang et al., 2022), which is also applied in the subsequent refinement process. Specifically, we obtain multiple judgments from the refiner and determine the final label through majority voting, as detailed in Appendix E.4. After majority voting, we randomly select one judgment that matches the voted label to serve as the final judgment. This process allows us to identify challenging prompts that elicit responses that do not accurately follow the instructions, yielding tuples in the form of $(x,y_{\text{negative}},j)$, where $y_{\text{negative}}$ is the incorrect response and $j$ is its corresponding judgment.
@@ 3.1 Experiment Setup
In this work, we focus on enhancing the instruction-following abilities of LLMs in a self-play fashion. We utilize SFT to bootstrap models under 10B parameters as actor and refiner models. For the more advanced LLaMA3-70B-Instruct, we directly employ it in both roles. Following this, we perform a three-iteration self-play training using 10k prompts per iteration from our generated dataset. In each iteration, we apply DPO for the actor and RFT for the refiner. We refer to the trained LLaMA3-8B-Instruct as SPaR-8B, LLaMA3-70B-Instruct as SPaR-70B, GLM-4-9B-Chat as SPaR-9B, and Mistral-7B-Instruct as SPaR-7B. More implementation details can be found in Appendix C. Description of baseline methods is provided in Appendix D.
@@ 3.3 Actor Evaluation Results
As shown in Appendix E.2, we assessed each iteration’s performance on general benchmarks, including GSM8k (Cobbe et al., 2021), TriviaQA (Joshi et al., 2017), MMLU (Hendrycks et al., 2020), and HumanEval (Chen et al., 2021). The results indicate that SPaR maintains or even improves general performance, particularly on GSM8k and HumanEval benchmarks, demonstrating that enhanced instruction-following capabilities support overall LLM alignment.
SPaR outperforms other baselines significantly.
Figure 3 demonstrates the improvements on IFEval with each training iteration. In every iteration, SPaR outperforms other methods. Notably,
@@
even after three iterations, other methods fail to surpass the performance of SPaR’s first iteration. Generally, our method and SELF outperform self-rewarding and meta-rewarding approaches, underscoring the importance of learning from refinement and excluding the interfering factors in instruction-following tasks. Furthermore, SPaR’s superior performance compared to SELF indicates that contrastive refinement response pairs can highlight key differences, which are difficult to learn using only correct responses. Additionally, only SPaR-8B-SFT outperforms the original LLaMA3-8B-Instruct, which suggests that incorporating the judgment SFT or refinement SFT data would reduce performance, likely due to the huge task gap and reduced diversity in the data.
@@ 3.5 Ablations and Analysis
Figure 5: Comparison of decoding strategies.
@@ 5 Conclusion
6 Acknowledgement
This work was supported by the National Science Foundation for Distinguished Young Scholars (with No. 62125604). This work was also supported by Tsinghua University Initiative Scientific Research Program. We would also like to thank Zhipu AI for sponsoring GPU computing and API cost consumed in this study.
@@ Appendix C Implementation Details
Appendix D Baselines
We compare our method with four popular self-improvement approaches, including:
Appendix E Experiment Results
E.1 Instruction-Following Evaluation Results.
@@ Appendix E Experiment Results
E.2 General Performance Evaluation
@@
E.3 Judgment Evaluation Results.
@@
E.4 Ablation Study on Judgment Capability.
@@
E.5 Ablation Study on Refinement Capability.
@@
E.6 Inference-time Scaling Comparison